OpenStudy (anonymous):

Identify function f(x),h(x) and g(x) such that the given function equals [f(g.h)](x)? a) ln (x^2 + 1)^1/2

OpenStudy (anonymous):

Is it f{[g(h)]}(x)? what does [f(g.h)](x) mean?

OpenStudy (anonymous):

yup2..f{[g(h)]}(x) :D

OpenStudy (anonymous):

if question a is ln[(x^2+1)^1/2], then, f(x)=ln(x), g(x)=x^1/2, h(x)=x^2+1. if question a is [ln(x^2+1)]^1/2, then, f(x)=x^1/2, g(x)=ln(x), h(x)=x^2+1.

OpenStudy (anonymous):

ouhhh..depend on us ryte?